OpenStudy (anonymous):

Please help, I have to have this in in 30 minutes. Joe has a collection of nickles and dimes that is worth $7.65. If the number of dimes was doubled and the number of nickles was increased by 35, the value of the coins would be $14.30. How many nickles and dimes does he have? Show your work.

OpenStudy (anonymous):

two equations: 10d+5n=765 10(2d)+5(n+35)=1430 solve for n and d: 5n=765-10d n=153-2d plug in to other equation: 20d+5((153-2d)+35)=1430 20d-10d+940=1430 10d=490 d=49 so 49 dimes and: n=153-2(49) n=55 49 dimes and 55 nickles